OpenStudy (anonymous):

What effect would you see if a wire was moved parallel to magnetic field lines?

OpenStudy (anonymous):

I think you will see no effect because magnetic field acts on a wire only when the angle between magnetic field vector B and the wire is not zero or in general Pi*n where n = 0,1,2,...

OpenStudy (anonymous):

agreed.... laurentz force will be generated only on perpendicular to the magnetic field....

OpenStudy (anonymous):

But I don't agree) Lorentz's force will act on a wire at any angle except Pi*n.

OpenStudy (anonymous):

F \[= q \left( v \times B \right)\] \[= I \left( L \times B \right)\]

OpenStudy (anonymous):

u sud agree ;)

OpenStudy (anonymous):

F = q v*B Sin Alpha Where alpha is the angle between vectors v and B.

OpenStudy (anonymous):

yes.... and alpha decides wht the angel is... right? alpha gives zero on pi*n....

OpenStudy (anonymous):

\[\sin \alpha \] .. sorry

OpenStudy (anonymous):

Yes right, but I said that Lorentz force exists at any angle EXCEPT Pi*n.